|
AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
|
CARME |
 |
|
Message ID: 508
Entry time: Mon Feb 12 01:40:48 2024
|
Author: |
TD, JM, CB, AR, RSS |
Subject: |
Run17 - Run20 |
|
|
02.35 DAQ continues file G22-88/R17_30
02.38 Check ASIC load
aida01 asics # 1 and #4 & aida12 asic #2 HEC events removed
09.05 DAQ continues file G22-88/R17_53
DSSSD bias & leakage current OK - attachments 10-11
FEE64 temperatures OK - attachment 12
ADC data item stats OK - attachment 13
hot channels aida01 ( 67k ) and aida15 ( 7k )
WR timestamps OK - attachment 14
per FEE64 Rate and Stat spectra - attachments 15-16
Merger, Merger stats and TapeServer - attachments 17-19
09.19 New run G22-88/R18
all histograms & stats zero'd
CB starting up 4He target preparatory to returning 16O2+ beam to CRYRING
L & R arms all out
09.26 4He target c. 2e11c.2
16O2+ beam c. 3e5/s
09.58 4He target c. 3e11/cm2
16O2+ beam c. 5e5/s at start of cycle
See attachments 20-22
10.40 New run G22-88/R19
all histograms & stats zero'd
CB will attempt to steer the beam
all steering options set to 0mm except YR09 target offset (vertical = -9mm) and (horizontal = 9mm)
10.52 R arm fully in, L arm fully out
beam steering hard right
14:24 Frank has informed us that the CRYRING team will be examining the electron cooler for th enext few hours. Run 19 stopped, Run 20 started.
Detectors moved fully out, target OFF.
All histograms and stats zero'd
Access to the cave is required, beam OFF
15:45 Channel 1 back to 150V.
17.40 DAQ continues file G22-88/R20_13
DSSSD bias & leakage current OK - attachments 23-24
FEE64 temperatures OK - attachment 25
All system wide checks OK except aida08 FPGA error - attachment 26
WR timestamps OK - attachment 27
ADC data item stats OK - attachment 28
hot channels aida01 ( 80k ) and aida15 ( 7k )
per FEE64 Rate and Stat spectra - attachments 29-30
Merger, Merger stats and TapeServer - attachments 31-32
23.12 DAQ continues file G22-88/R20_20
DSSSD bias & leakage current OK - attachments 33-34
FEE64 temperatures OK
All system wide checks OK except aida08 FPGA error - attachment 35
WR timestamps OK - attachment 36
ADC data item stats OK - attachment 37
hot channels aida01 ( 80k ) and aida15 ( 7k )
per FEE64 Rate and Stat spectra - attachments 38
Merger, Merger stats and TapeServer - attachments 39-40
03.47 DAQ continues file G22-88/R20_28
DSSSD bias & leakage current OK - attachments 41-42
FEE64 temperatures OK - attachment 43
ADC data item stats - attachment 44
aida12 rate - HEC events
per FEE64 Rate spectra - attachment 45
Check ASIC load
ADC data item stats - attachment 46
per FEE64 Rate spectra - attachment 47
Merger, Merger stats and TapeServer - attachments 48-49
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|
|